# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ Abstract
Altermagnets is a class of antiferromagnetic materials which has electron bands with lifted spin degeneracy in momentum space but vanishing net magnetization and no stray magnetic fields. Because of these properties, altermagnets have attracted much attention for potential use in spintronics. We here show that despite the absence of bulk magnetization, the itinerant electrons in altermagnets can generate a magnetization close to edges and vacuum interfaces. We find that surface-induced magnetization can also occur for conventional antiferromagnets with spin-degenerate bands, where the magnetization from the itinerant electrons originates from a subtle yet nonvanishing redistribution of the probability density on the unit cell level. An intuitive explanation of this effect in a phenomenological model is provided. In the altermagnetic case, the induced magnetization has a different spatial dependence than in the antiferromagnetic case due to the anisotropy of the spin-polarized Fermi surfaces, causing the edge-induced Friedel oscillations of the spin-up and spin-down electron densities to have different periods. We employ both a low-energy effective continuum model and lattice tight-binding calculations. Our results have implications for the usage of altermagnets and antiferromagnets in nanoscale spintronic applications.
@@ I Introduction
The spin polarization of quasiparticles in solid-state materials is a key property for many quantum physical phenomena, not the least in the field of spintronics. Conventional antiferromagnetic materials feature spin-degenerate electron bands, as dictated by parity-time-reversal ($\mathcal{PT}$) symmetry. However, it has recently been understood that antiferromagnetic materials can display a momentum-dependent spin-splitting which is distinct from relativistically spin-orbit coupled systems. One such mechanism, originally envisioned in Ref. [1], consists of a spin-lattice coupling due to an internal periodic magnetic field in the material. When $\mathcal{PT}$ symmetry is broken by the combined lattice geometry and spin ordering in a material, a large spin splitting can arise which depends on momentum $\boldsymbol{k}$ [2, 3, 4, 5, 6, 7], while maintaining zero net magnetization upon integration over the entire first Brillouin zone  [^1] According to the extended classification in Ref. [8], altermagnets with vanishing magnetization belong to types II and III altermagnets. Such materials have been named altermagnets [9, 10], and ab initio calculations have identified several possible material candidates that can host an altermagnetic state [11]. This includes metals like RuO2 [3, 7, 12] and Mn5Si3 [13] as well as semiconductors/insulators like MnTe [12, 9], CrSb [9], MnF2 [4, 14], and La2CuO4 [9]. Recent ARPES measurements in MnTe [15, 16, 17], RuO2 [18, 19], and CrSb [20, 21, 22] have corroborated several of these predictions  [^2] The magnetic order in RuO2 is debated [23, 24].. Another piece of evidence supporting altermagnetism is provided by measuring the anomalous Hall effect [25, 26] and spin-splitter effect and torques [27, 28, 29]. It is noteworthy that a Fermi-surface instability caused by quasiparticle interactions has been predicted to cause an altermagnetic spin-splitting in the band structure [30] long before the recent surge of activity in the field. On the other hand, such an instability is interestingly restricted for $p$-wave magnetism [31].
@@
In this work, we answer this question and show that altermagnets generically develop a nonvanishing magnetization at interfaces, except for very specific crystallographic orientations of the interface. This effect originates from the explicit breakdown of the symmetry between the spin-polarized Fermi surfaces of altermagnets. Then, the periods of the Friedel oscillations of the spin-up and spin-down electron densities become different leading to oscillating magnetization near the edges. The maximal breakdown is achieved if one of the Fermi surfaces is elongated along the confinement direction, i.e., $\varepsilon_{\uparrow,\mathbf{k}_{\parallel}}\neq\varepsilon_{\downarrow,-\mathbf{k}_{\parallel}}$, where $\mathbf{k}_{\parallel}$ is the momentum component along the interface and $\varepsilon_{\sigma,\mathbf{k}}$ denotes the dispersion relation for spin $\sigma$. The symmetry between the Fermi surfaces and, as a result, the vanishing magnetization is preserved only for a specific orientation of the lobes such that $\varepsilon_{\uparrow,\mathbf{k}_{\parallel}}=\varepsilon_{\downarrow,-\mathbf{k}_{\parallel}}$. For any other orientation, a magnetization develops near the interface of the altermagnet, which means that it is a robust phenomenon.
@@
In our work, we combine analysis via a low-energy effective continuum model with a tight-binding lattice calculation to address the contribution of itinerant electrons in the magnetization of metallic altermagnets and antiferromagnets. The results of the lattice calculations are more nuanced and contain information on the structure of the wave functions on the unit cell level. While such resolution is not necessary to obtain an interface-induced magnetization in altermagnets, which has a strong contribution from the spin-split Fermi surface, it allows us to analyze the edge magnetization also in antiferromagnets and show that it too may demonstrate an oscillating behavior despite the spin-degenerate electron bands in the bulk. The latter originates from the redistribution of the electron wave functions near the uncompensated interfaces and is distinct from the edge magnetization in altermagnets which also contains the contribution from the spin-dependent Fermi momentum. The appearance of magnetization in antiferromagnets and altermagnets has important implications with regard to their potential role in nanoscale spintronic applications.
@@
The focus of our work on itinerant electrons distinguishes it from other very recent works on surface magnetization that was studied primarily in insulating antiferromagnets [38, 39, 40, 41, 42, 43], where the main attention is devoted to the distortion of the localized spin moment texture of the lattice sites near the surface. Our model setup and results are also distinct from the recent proposal to realize magnetization in bent altermagnets [44], where the Néel order parameter becomes nonuniform.
@@
Our paper is organized as follows. In Sec. II, we introduce an effective low-energy model and provide an analytical calculation of the local density of state (LDOS) and magnetization in an altermagnetic ribbon; an antiferromagnetic ribbon is also discussed. Characterization of the interface-induced magnetization in a tight-binding model is provided in Sec. III. The results are discussed and summarized in Sec. IV. Throughout this paper, we use $\hbar=k_{\rm B}=1$.
@@ II.1 Antiferromagnets
As a warm-up, we consider an antiferromagnetic ribbon, see Fig. 1 for the lattice geometry. Since an antiferromagnetic material has spin-degenerate energy bands, we expect no magnetization arising due to itinerant electrons unless a spin asymmetry is introduced via, e.g., boundary conditions. To illustrate this, in what follows, we consider a ribbon of an antiferromagnetic metal, and show that the corresponding LDOS remains spin-degenerate, i.e., no magnetization arises in a continuum model. However, this changes in the tight-binding model where the unit cell geometry gives rise to phase shifts of the itinerant electron wave functions which turn out to produce a net interface magnetization, as we discuss later in Sec. III.3. This suggests that the standard boundary conditions of a vanishing wave function at the vacuum edge are insufficient to capture the interfacial physics in antiferromagnetic systems.
@@ II.1.1 Model and eigenfunctions in antiferromagnetic ribbon
The low-energy continuum Hamiltonian for an antiferromagnetic material with two spin-full sites in the unit cell is
$H_{\rm AF}(\mathbf{k})=\begin{pmatrix}-sJ_{sd}-\mu&\varepsilon_{0}-\frac{k^{2}}{2m}\\ \varepsilon_{0}-\frac{k^{2}}{2m}&sJ_{sd}-\mu\end{pmatrix},$ (1)
where $\mu$ is the chemical potential, $m$ is the effective mass, $\mathbf{k}$ is the momentum, $J_{sd}$ is the strength of the $s-d$ coupling, $s=\pm$ is the spin projection, and $\varepsilon_{0}$ has the dimension of energy. In what follows, we find it convenient to introduce dimensionless variables: $\tilde{\varepsilon}=\varepsilon/\mu$, $\tilde{k}=k/k_{F}$ with $k_{F}=\sqrt{2m\mu}$, and $\tilde{J}=J_{sd}/\mu$.
The energy spectrum of the model (1) reads
$\tilde{\varepsilon}_{\pm}=-1\pm\sqrt{\tilde{J}^{2}+\left(\tilde{\varepsilon}_{0}-\tilde{k}^{2}\right)^{2}}$ (2)
and describes two spin-degenerate bands separated by a gap.
For simplicity, we consider the case where only a lower band $\tilde{\varepsilon}_{-}$ is filled. Then, by projecting out the higher energy band and expanding the obtained Hamiltonian up to the leading nontrivial order in $\tilde{k}$, we obtain:
$\tilde{H}_{\rm AF,-}(\mathbf{k})=-\left(1+\sqrt{\tilde{J}^{2}+\tilde{\varepsilon}_{0}^{2}}\right)+\frac{\tilde{\varepsilon}_{0}}{\sqrt{\tilde{J}^{2}+\tilde{\varepsilon}_{0}^{2}}}\tilde{k}^{2},$ (3)
which can be further simplified by introducing a new set of dimensionless variables: $\varepsilon^{\prime}=\varepsilon/\left[\mu\left(1+\sqrt{\tilde{J}^{2}+\tilde{\varepsilon}_{0}^{2}}\right)\right]$ and $k^{\prime}=k/q_{F}$ with $q_{F}=\sqrt{2m\mu\left(1+\sqrt{\tilde{J}^{2}+\tilde{\varepsilon}_{0}^{2}}\right)\sqrt{\tilde{J}^{2}+\tilde{\varepsilon}_{0}^{2}}/\tilde{\varepsilon}_{0}}$.
In the ribbon geometry shown in Fig. 1, only the momentum along the edges, $k_{y}$, remains a good quantum number. For the sake of definiteness, we direct the $x$-axis normal to the ribbon and the $y$-axis along its edge. The antiferromagnetic ribbon is described by the Hamiltonian (3) with $k_{x}\to-i\nabla_{x}$.
@@
Figure 1: The antiferromagnetic (AFM) and altermagnetic (AM) ribbons are modelled both with a continuum model and a tight-binding lattice model with a four-site unit cell which allows us to continuously tune the system from AFM to AM. In the lattice model, sites 1 and 4 are localized, anti-collinear magnetic moments, coupled to the itinerant electrons through an s-d coupling. Sites 2 and 3 are non-magnetic sites with an on-site energy $\varepsilon_{i}$. The unit cell indices are $(I,J)$, while the site indices are $(i,j)$. To model an AFM, we take $\varepsilon_{2}=\varepsilon_{3}=0$ while $\varepsilon_{2}\neq\varepsilon_{3}$ gives rise to AM order by introducing anisotropic spin-splitting mediated by the non-magnetic sites. The nearest-neighbor hopping is described by $t$.
@@ II.1.2 Density of states and magnetization
$\nu_{s}(\varepsilon,x)=\sum_{n=0}^{\infty}\int\frac{dk_{y}}{2\pi}|\psi_{s,k_{y},n}(x)|^{2}\,\delta{\left(\varepsilon-\varepsilon_{n}\right)},$ (9)
@@
where $\nu_{0}=q_{F}^{2}/\left[4\pi\mu\left(1+\sqrt{\tilde{J}^{2}+\tilde{\varepsilon}_{0}^{2}}\right)\right]$ is the bulk DOS. As is evident from Eq. (10), the LDOS in an antiferromagnetic ribbon is spin-independent. Therefore, there is no magnetization in this model. To set up the stage for the comparison with altermagnets in Sec. II.2, we calculate the LDOS and the local charge density in a few approximations.
@@
$\lim_{x^{\prime}\to\infty}\!\nu_{s}(\varepsilon^{\prime},x^{\prime})$$\approx$ $\nu_{0}\!\left[1-\frac{1}{\sqrt{\pi}\sqrt{x^{\prime}}}\frac{\sin{\left(2\sqrt{\varepsilon^{\prime}+1}x^{\prime}+\pi/4\right)}}{\left(\varepsilon^{\prime}+1\right)^{1/4}}\right],$
$\lim_{x^{\prime}\to\infty}\!n_{s}(x^{\prime})$$\approx$ $n_{0}\!\left[1+\frac{1}{\sqrt{\pi}x^{\prime 3/2}}\cos{\left(2x^{\prime}+\frac{\pi}{4}\right)}\right].$ (15)
@@
Thus, since the energy bands in the AFM are spin-degenerate and the BCs in the continuum model do not distinguish spin projections, the LDOS and the charge density in the antiferromagnets remain spin-degenerate. Neither local nor net magnetizations occur in this case. As mentioned at the beginning of this section, we will see that this conclusion changes in a tight-binding model discussed in Sec. III.

### v3

@@ Abstract
Altermagnets is a class of antiferromagnetic materials which has electron bands with lifted spin degeneracy in momentum space but vanishing net magnetization and no stray magnetic fields. Because of these properties, altermagnets have attracted much attention for potential use in spintronics. We here show that despite the absence of bulk magnetization, the itinerant electrons in altermagnets can generate a magnetization close to edges and vacuum interfaces. We find that surface-induced magnetization can also occur for conventional antiferromagnets with spin-degenerate bands, where the magnetization from the itinerant electrons originates from a subtle yet nonvanishing redistribution of the probability density on the unit-cell level. An intuitive explanation of this effect in a phenomenological model is provided. In the altermagnetic case, the induced magnetization has a different spatial dependence than in the antiferromagnetic case due to the anisotropy of the spin-polarized Fermi surfaces, causing the edge-induced Friedel oscillations of the spin-up and -down electron densities to have different periods. We employ both a low-energy effective continuum model and lattice tight-binding calculations. Our results have implications for the usage of altermagnets and antiferromagnets in nanoscale spintronic applications.
@@ I Introduction
The spin polarization of quasiparticles in solid-state materials is a key property for many quantum physical phenomena, not the least in the field of spintronics. Conventional antiferromagnetic materials feature spin-degenerate electron bands, as dictated by parity-time-reversal ($\mathcal{PT}$) symmetry. However, it has recently been understood that antiferromagnetic materials can display a momentum-dependent spin-splitting which is distinct from relativistically spin-orbit coupled systems. One such mechanism, originally envisioned in Ref. [1], consists of a spin-lattice coupling due to an internal periodic magnetic field in the material. When $\mathcal{PT}$ symmetry is broken by the combined lattice geometry and spin ordering in a material, a large spin splitting can arise which depends on momentum $\boldsymbol{k}$ [2, 3, 4, 5, 6, 7], while maintaining zero net magnetization upon integration over the entire first Brillouin zone.  [^1] According to the extended classification in Ref. [8], altermagnets with vanishing magnetization belong to types II and III altermagnets. Such materials have been named altermagnets [9, 10], and ab initio calculations have identified several possible material candidates that can host an altermagnetic state [11]. This includes metals like RuO2 [3, 7, 12] and Mn5Si3 [13] as well as semiconductors/insulators like MnTe [12, 9], CrSb [9], MnF2 [4, 14], and La2CuO4 [9]. Recent ARPES measurements in MnTe [15, 16, 17], RuO2 [18, 19], and CrSb [20, 21, 22] have corroborated several of these predictions  [^2] The magnetic order in RuO2 is debated [23, 24].. Another piece of evidence supporting altermagnetism is provided by measuring the anomalous Hall effect [25, 26] and spin-splitter effect and torques [27, 28, 29]. It is noteworthy that a Fermi-surface instability caused by quasiparticle interactions has been predicted to cause an altermagnetic spin-splitting in the band structure [30] long before the recent surge of activity in the field. On the other hand, such an instability is interestingly restricted for $p$-wave magnetism [31].
@@
In this work, we answer this question and show that altermagnets generically develop a nonvanishing magnetization at interfaces, except for very specific crystallographic orientations of the interface. This effect originates from the explicit breakdown of the symmetry between the spin-polarized Fermi surfaces of altermagnets. Then, the periods of the Friedel oscillations of the spin-up and -down electron densities become different leading to oscillating magnetization near the edges. The maximal breakdown is achieved if one of the spin-polarized Fermi surfaces is elongated along the confinement direction, i.e., $\varepsilon_{\uparrow,\mathbf{k}_{\parallel}}\neq\varepsilon_{\downarrow,-\mathbf{k}_{\parallel}}$, where $\mathbf{k}_{\parallel}$ is the momentum component along the interface and $\varepsilon_{\sigma,\mathbf{k}}$ denotes the dispersion relation for spin $\sigma$. The symmetry between the Fermi surfaces and, as a result, the vanishing magnetization is preserved only for a specific orientation of the lobes such that $\varepsilon_{\uparrow,\mathbf{k}_{\parallel}}=\varepsilon_{\downarrow,-\mathbf{k}_{\parallel}}$. For any other orientation, a magnetization develops near the interface of the altermagnet, which means that it is a robust phenomenon.
@@
In our work, we combine analysis via a low-energy effective continuum model with a tight-binding lattice calculation to address the contribution of itinerant electrons in the magnetization of metallic altermagnets and antiferromagnets. The results of the lattice calculations are more nuanced and contain information on the structure of the wave functions on the unit-cell level. While such resolution is not necessary to obtain an interface-induced magnetization in altermagnets, which has a strong contribution from the spin-split Fermi surface, it allows us to analyze the edge magnetization also in antiferromagnets and show that it too may demonstrate an oscillating behavior despite the spin-degenerate electron bands in the bulk. The latter originates from the redistribution of the electron wave functions near the uncompensated interfaces and is distinct from the edge magnetization in altermagnets which also contains the contribution from the spin-dependent Fermi momentum. The appearance of magnetization in antiferromagnets and altermagnets has important implications with regard to their potential role in nanoscale spintronic applications.
@@
The focus of our work on itinerant electrons distinguishes it from other very recent works on surface magnetization that was studied primarily in insulating antiferromagnets [38, 39, 40, 41, 42, 43]. While the main attention in these works is devoted to the distortion of the localized spin moment texture of the lattice sites near the surface, in our work, we ignore any redistribution and dynamics of magnetic atoms. This allows us to distinguish the edge magnetization stemming from a different origin and identify universal features that are determined by the bulk properties rather than details of the interface. The treatment of the spins of magnetic atoms as a static background is justified in the limit of low temperatures, which we assume in our work. Our model setup and results are also distinct from the recent proposal to realize magnetization in bent altermagnets [44], where the Néel order parameter becomes nonuniform.
@@
Our paper is organized as follows. In Sec. II, we introduce an effective low-energy model and provide an analytical calculation of the local density of state (LDOS) and magnetization in an altermagnetic ribbon; an antiferromagnetic ribbon is also discussed. Characterization of the interface-induced magnetization in a tight-binding model is provided in Sec. III. Technical details related to the derivation of the continuum models and the definition of the spectral function in the lattice model are presented in Appendixes A and B, respectively. The results are discussed and summarized in Sec. IV. Throughout this paper, we use $\hbar=k_{\rm B}=1$.
@@ II Effective continuum model: theory and results
We start our investigation of edge-induced magnetization from the analysis in continuum low-energy models. Such models are derived from the lattice model, see Sec. III and Appendix A, and are intended to capture the key aspects of materials. As we demonstrate in our work, while being successful in predicting the bulk properties and altermagnetic surface effects, the inherent coarse-graining of continuum models does not allow to capture a subtle interplay between boundaries and the structure of the unit-cell. Nevertheless, the use of continuum models is instructive and allows us to separate different mechanisms responsible for edge-induced magnetization.
@@ II.1 Antiferromagnets
As a warm-up, we consider an antiferromagnetic ribbon, see Fig. 1(a) for the setup geometry. Since an antiferromagnetic material has spin-degenerate energy bands, we expect no magnetization arising due to itinerant electrons unless a spin asymmetry is introduced via, e.g., boundary conditions. To illustrate this, in what follows, we consider a ribbon of an antiferromagnetic metal, and show that the corresponding LDOS remains spin-degenerate, i.e., no magnetization arises in a continuum model. However, this changes in the tight-binding model where the unit-cell geometry gives rise to phase shifts of the itinerant electron wave functions which turn out to produce a net interface magnetization, as we discuss later in Sec. III.3. This suggests that the standard boundary conditions of a vanishing wave function at the vacuum edge are insufficient to capture the interfacial physics in antiferromagnetic systems.
@@ II.1.1 Model and eigenfunctions in antiferromagnetic ribbon
The simplest low-energy continuum Hamiltonian for an antiferromagnet is
$H_{\rm AFM}(\mathbf{k})=\frac{\varepsilon_{0}}{\sqrt{J_{\rm sd}^{2}+\varepsilon_{0}^{2}}}\frac{k^{2}}{2m}-\left(\mu+\sqrt{J_{\rm sd}^{2}+\varepsilon_{0}^{2}}\right),$ (1)
where $\mu$ is the chemical potential, $m$ is the effective mass, $\mathbf{k}$ is the momentum, $J_{\rm sd}$ is the strength of the s-d coupling, and $\varepsilon_{0}$ is a constant with the dimension of energy. In what follows, we find it convenient to introduce dimensionless variables: $\tilde{\varepsilon}=\varepsilon/\mu$, $\tilde{k}=k/k_{F}$ with $k_{F}=\sqrt{2m\mu}$, and $\tilde{J}=J_{\rm sd}/\mu$. The Hamiltonian (1) does not depend on the spin projection $s=\pm$ and describes two spin-degenerate energy bands.
The low-energy continuum model (1) can be straightforwardly derived from the lattice model defined in Sec. III.1 if one retains only the lowest-energy band and performs an expansion in momentum around the $\Gamma$ point, see Appendix A for details of the derivation. Being a continuum model, the model (1) loses information about the unit-cell structure. In particular, it becomes insensitive to features defined on the scale of the lattice constant where a more refined lattice treatment is required. Yet, as we will phenomenological show in Sec. III.4, the continuum model can still be useful in providing a qualitative picture of the lattice results.
The Hamiltonian (1) can be further simplified by introducing a new set of dimensionless variables: $\varepsilon^{\prime}=\varepsilon/\left[\mu\left(1+\sqrt{\tilde{J}^{2}+\tilde{\varepsilon}_{0}^{2}}\right)\right]$ and $k^{\prime}=k/q_{F}$ with $q_{F}=\sqrt{2m\mu\left(1+\sqrt{\tilde{J}^{2}+\tilde{\varepsilon}_{0}^{2}}\right)\sqrt{\tilde{J}^{2}+\tilde{\varepsilon}_{0}^{2}}/\tilde{\varepsilon}_{0}}$. Then, the dimensionless version of the Hamiltonian (1) reads
$H_{\rm AFM}(\mathbf{q})=\mu\left(1+\sqrt{\tilde{J}^{2}+\tilde{\varepsilon}_{0}^{2}}\right)\left(\tilde{q}^{2}-1\right).$ (2)
In the ribbon geometry shown in Fig. 1(a), only the momentum along the edges remains a good quantum number. For the sake of definiteness, we direct the $x$ axis normal to the ribbon and the $y$ axis along its edge. The antiferromagnetic ribbon is described by the Hamiltonian (2) with $k_{x}\to-i\nabla_{x}$.
@@
Figure 1: The antiferromagnetic (AFM) and altermagnetic (AM) ribbons are modeled both with a continuum model (a) and a tight-binding lattice model (b) with a four-site unit-cell which allows us to continuously tune the system from AFM to AM. In the continuum model, the ribbon is infinite in the y-direction and of length L in the x-direction. In the lattice model, sites 1 and 4 are localized, anti-collinear magnetic moments, coupled to the itinerant electrons through an s-d coupling. Sites 2 and 3 are nonmagnetic sites with an on-site energy $\varepsilon_{i}$. The unit-cell indices are $(I,J)$, while the site indices are $(i,j)$. To model an AFM, we take $\varepsilon_{2}=\varepsilon_{3}=0$ while $\varepsilon_{2}\neq\varepsilon_{3}$ gives rise to AM order by introducing anisotropic spin-splitting mediated by the nonmagnetic sites. The nearest-neighbor hopping is described by $t$.
@@ II.1.2 Density of states and magnetization
$\nu_{s}(\varepsilon,x)=\sum_{n=0}^{\infty}\int\frac{dk_{y}}{2\pi}|\psi_{s,k_{y},n}(x)|^{2}\,\delta{\left(\varepsilon-\varepsilon_{n}\right)}.$ (8)
@@
where $\nu_{0}=q_{F}^{2}/\left[4\pi\mu\left(1+\sqrt{\tilde{J}^{2}+\tilde{\varepsilon}_{0}^{2}}\right)\right]$ is the bulk DOS. As is evident from Eq. (9), the LDOS in an antiferromagnetic ribbon is spin independent. Therefore, there is no edge-induced magnetization in this continuum model. To set up the stage for the comparison with altermagnets in Sec. II.2, we calculate the LDOS and the local charge density in a few approximations.
@@
$\nu_{s}(\varepsilon^{\prime},x^{\prime})$ $=$ $\nu_{0}\!\left[1-\frac{1}{\sqrt{\pi x^{\prime}}}\frac{\sin{\left(2\sqrt{\varepsilon^{\prime}+1}x^{\prime}+\pi/4\right)}}{\left(\varepsilon^{\prime}+1\right)^{1/4}}\right]+\mathcal{O}{\left(\frac{1}{x^{\prime}}\right)},$
$n_{s}(x^{\prime})$ $=$ $n_{0}\!\left[1+\frac{1}{\sqrt{\pi}(x^{\prime})^{3/2}}\cos{\left(2x^{\prime}+\frac{\pi}{4}\right)}\right]+\mathcal{O}{\left(\frac{1}{(x^{\prime})^{2}}\right)}.$
@@
Thus, since the energy bands in the AFM are spin-degenerate and the boundary conditions in the continuum model do not distinguish spin projections, the LDOS and the charge density in the continuum model of antiferromagnets remain spin-degenerate. Neither local nor net magnetizations occur in this case. As mentioned at the beginning of this section, we will see that this conclusion changes in a tight-binding model discussed in Sec. III.